User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!


Katherine 12m:

Disk VSN: HOB+0124

Data volume at beginning: 3396GB

  • 1700UT - Good start, everything peachy (JS)
  • 2200UT - Getting some DBBC offset drifts (JS)
  • 2205UT - Reset DBBC during gap, still missed scans 159-2210,159-2212a (JS)
  • 0030UT - Heater switched off in maser room (Approximate time given by Bret, may have been >0033UT) (JR)
  • 0033UT - ALARM: Delay through DBBC has changed. Will monitor and reset DBBC if problem persists. (JR)
  • 0825UT - ALARM: Delay through DBBC has changed. Monitoring for now (Tiege)
  • 0836UT - Still drifting, Have halted the schedule and attempting to reset dbbc and fix the problem. Scans 160-0825, 160-0828a, 160-0834 were all undertaken during the drifting. (Tiege)
  • 0908UT - Think I have fixed the issue, first two attempts at stopping the drifting were unsuccessful, drifting occurred again within minutes of resetting DBBC, additionally mk5 spat the dummy and dimino needed to be killed and restarted. Scans missed were 160-0836a through to 160-0910b inclusive. First good scan is 160-0918a. (Tiege)
  • 2016.160.13:04:28 ALARM: Delay through DBBC has changed. Monitor for stability and reset DBBC if it drifts. (Lucas) Only one alarm so far but will reconfigure if it persists.

Yarragadee 12m:

Disk VSN: HOB+0128

Data volume at beginning: 2067GB

  • 1700UT - Good start, everything peachy (JS)
/home/www/auscope/opswiki/data/pages/handover/r1743.txt · Last modified: 2016/06/08 13:07 by Jesse Swan